The solutions to the inequality (2x+3)1/2 >lxl exist within which interval?

A. the interval -3/2<x<-1

B. the interval -x>3/2

C. the interval -3/2<x<3


1
Expert's answer
2021-08-09T16:53:03-0400
"2x+3\geq0=>x\geq-\dfrac{3}{2}"

"2x+3>x^2"

"x^2-2x-3<0"

"(x+1)(x-3)<0"

"-1<x<3, x\geq-\dfrac{3}{2}"

The solutions are "-1<x<3."

Interval "(-1,3)" lies within the interval "-\dfrac{3}{2}<x<3."